Find the amount which Ram will get on โ‚น 4,096, he gave it for 18 months at 1๐Ÿ whole ๐Ÿ/๐Ÿ per annum, interest being compounded half yearly.ย 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

P =โ‚น4,096
R = 12 whole ½ per annum =25/2 per annum=25/4 per half year
n = 18 months
There will be 3 half years in 18 months. Therefore,
Amount, A =P (1 +R/100 )^n
= โ‚น [4096(1+25/4 x 100 )^3]
= โ‚น [4096×(1+1/16)^3]
= โ‚น 4096×(17/16)^3
= โ‚น 4913
∴ The required amount isโ‚น4,913.

โ€‹Thanks